OpenStudy (calculusxy):

Maya chatted with all her friends for 2/3 of an hour on Saturday and 1/4 of an hour on Sunday. During that time period, she chatted with Paul for 1/8 of an hour. What fraction of the time Maya spent chatting with her friends online was spent chatting with Paul?

OpenStudy (calculusxy):

@wio

OpenStudy (anonymous):

Find the total time she spent chatting.

OpenStudy (calculusxy):

I got 11/96 because I added 8/12+3/12=11/12 => 11/12*1/8=11/96. Am I correct?

OpenStudy (anonymous):

Find the time she spend chatting with Paul. Divide this by the time spent chatting.

OpenStudy (calculusxy):

Am i correct?

OpenStudy (calculusxy):

@agent0smith

OpenStudy (calculusxy):

Doesn't "of" mean to multiply?

OpenStudy (anonymous):

yeah

OpenStudy (calculusxy):

So am i correct?

OpenStudy (anonymous):

I don't think so.

OpenStudy (calculusxy):

Well the choices that I got are: 11/96 3/22 19/24 25/24

OpenStudy (anonymous):

I think they want: \[ \frac{1/8}{2/3+ 1/4+1/8} \]

OpenStudy (anonymous):

You might be correct, I just don't know that you are doing it correctly.

OpenStudy (calculusxy):

I dont know why but while computing: \[\frac{ \frac{ 1 }{ 8 } }{ \frac{ 2 }{ 3 }+\frac{ 1 }{ 4 }+\frac{ 1 }{ 8 } } = \frac{ 3 }{ 25 }\] For some reason.

OpenStudy (anonymous):

oh wait! I misread. It should be \[ \frac{1/8}{2/3+1/4} \]

OpenStudy (calculusxy):

So you're saying that it should be 3/22? Can you explain how you got this equation?

OpenStudy (anonymous):

Well. \(2/3+1/4\) is the total.

OpenStudy (anonymous):

I did \[ \frac{chat\ with\ Paul}{chat\ total} \]

OpenStudy (calculusxy):

No pun intended, but doesnt "What fraction of the time.." mean multiplying?

OpenStudy (anonymous):

Yes, the "of" here means that we can multiply this fraction by the total time to get the time spent chatting will Paul.

OpenStudy (calculusxy):

So what do you think we should do?

OpenStudy (anonymous):

Basically we have: \[ \text{fraction} \times \text{chat total} = \text{chat with Paul} \]Then, it becomes: \[ \text{fraction} = \frac{ \text{chat with Paul}}{\text{chat total}} \]

OpenStudy (calculusxy):

What do you mean by "fractions"?

OpenStudy (anonymous):

So, it was multiplication, but we had to divide to get the part we wanted.

OpenStudy (calculusxy):

Sorry for asking all these questions, but why should we divide instead of multiply?

OpenStudy (calculusxy):

To get the amount of time Maya chatted with Paul.

OpenStudy (calculusxy):

I really need this to be correct. I might get graded.

OpenStudy (anonymous):

What fraction of the time Maya spent chatting with her friends online was spent chatting with Paul? \[ \text{What } \color{red}{\text{fraction}} \text{ of the } \color{blue}{\text{time Maya spend chatting with her friends online}} \\\\ \text{ was spent } \color{green}{\text{chatting with Paul}}\text{?} \] The red part is what we are looking for. This corresponds to \[ \color{red}{\text{fraction}} \times \color{blue}{\text{total chatting online}} = \color{green}{\text{time chatting with Paul}} \]Then, it becomes: \[ \color{red}{\text{fraction}} = \frac{ \color{green}{\text{time chatting with Paul}} }{ \color{blue}{\text{total chatting online}} } \]

OpenStudy (calculusxy):

SO you are saying that if I multiply 3/22 by 11/12, i will get 1/8?

OpenStudy (anonymous):

Yes.

OpenStudy (calculusxy):

Thanks.